# Learning Effective NeRFs and SDFs Representations with 3D Generative Adversarial Networks for 3D Object Generation: Technical Report for ICCV 2023 OmniObject3D Challenge

### v1

@@ Abstract
In this technical report, we present a solution for 3D object generation of ICCV 2023 OmniObject3D Challenge. In recent years, 3D object generation has made great process and achieved promising results, but it remains a challenging task due to the difficulty of generating complex, textured and high-fidelity results. To resolve this problem, we study learning effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Specifically, inspired by recent works, we use the efficient geometry-aware 3D GANs as the backbone incorporating with label embedding and color mapping, which enables to train the model on different taxonomies simultaneously. Then, through a decoder, we aggregate the resulting features to generate Neural Radiance Fields (NeRFs) based representations for rendering high-fidelity synthetic images. Meanwhile, we optimize Signed Distance Functions (SDFs) to effectively represent objects with 3D meshes. Besides, we observe that this model can be effectively trained with only a few images of each object from a variety of classes, instead of using a great number of images per object or training one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
@@
1 Introduction
To devise a solution for 3D object generation in the ICCV 2023 OmniObject3D Challenge, we simultaneously consider effective 3D representation learning and advanced generation strategy. In terms of 3D representation learning, textured mesh [3] and Neural Radiance Fields (NeRFs) [14] embrace the great potential to effectively encode 3D object features. In particular, textured mesh representations can support individual texture map generation, which allows us to easily replace surface textures, while NeRFs-based representations excel at rendering high-fidelity images. Furthermore, when considering effective shape representation, there are a variety of options, ranging from explicit representations like point-based [20], mesh-based [1] and voxel-based [22] approaches to implicit ones such as the Signed Distance Functions (SDFs) [17, 11] and occupancy function [13]. In the light of these cutting-edge techniques, we mainly focus on exploring effective NeRFs representations for 3D object generation while still optimizing SDFs representations to effectively represent objects with 3D meshes.
Moving on to advanced generation strategy, there are usually different options, such as latent generation [21] and direct generation [16]. In particular, latent generation methods aim to generate the representation of the desired output in the latent space and decode the latent code back to the 3D space, while direct generation focus on directly generating the desired 3D representation without utilizing the latent space. Besides, from the perspective of generation process, generation strategies can be categorized into Generative Adversarial Networks (GANs) [10, 2], Variational AutoEncoder (VAE) [19], flow-based [6], Diffusion Models (DMs) [21], etc. In our solution, we employ the efficient geometry-aware 3D GANs [2] as the backbone for 3D object generation.
@@ 1 Introduction
The overall framework of our solution is depicted in Fig. 1. Specifically, we aim to learn effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Inspired by the recent success of efficient geometry-aware 3D GANs, i.e., EG3D [2, 18], we employ it as the backbone and incorporate label embedding and color mapping [18] to enable model training on different taxonomies simultaneously. Next, we generate NeRFs-based representations for rendering high-fidelity synthetic images and SDFs to effectively represent objects with 3D meshes through a decoder. In addition, we empirically find that we can effectively train this model with only a few images per object from various classes, rather than training with a great number of images per object or optimizing one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
2 Technical Solution
As aforementioned, to realize 3D object generation for ICCV 2023 OmniObject3D Challenge, we explore 3D GANs to learn effective NeRFs representations for synthetic images and SDFs representations for 3D shapes.
Motivation of Using GANs instead of DMs.
@@ 2 Technical Solution
Model Training.
Since we are simultaneous optimizing NeRFs and SDFs representations, we ensure the generator can initially generate a unit sphere by initializing our model with SDF-pretraining as [18]. Then, the generator is optimized using the Adam optimizer with a learning rate of 0.0025, while the discriminator [2] uses a learning rate of 0.002. The batch size is set to 32 and the model is trained for 300,000 iterations. Additionally, we also use adaptive discriminator augmentation [8] to improve model generalization. Overall, the model training objective is formulated as:
@@
Figure 2: Visualization results of the synthesis images rendered from the NeRFs generated by our solution.
3 Experiments
Dataset.
@@ 3 Experiments
Efficient Model Training.
Instead of training one model for each class, we train a single model with various types of objects. The model takes each image along with its camera pose, focal length and object type. Besides, we empirically find that we can effectively train this model with only a few images per object from various classes, so we select the first 8 images of each object for model training. This significantly improves the model training efficiency.
Figure 3: Visualization results of the synthesis images rendered from the NeRFs generated by our solution without additional label embedding.
@@
$\Delta$= Ours w/ label embedding - -4.8
Ours w/o label embedding
Table 1: Improvement in terms of FID of our solution after using the additional label embedding.
Results.
Fig. 2 shows some visualization results of synthesis images generated by our solution on OmniObject3D. It can be seen that our approach is capable of generating high-fidelity images from different views for 3D object generation. Besides, Fig. 3 shows results of our solution without the additional label embedding. Comparing Fig. 2 with Fig. 3, we can see that without the additional label embedding, the quality of the synthesis images is significantly worse than the original solution. Furthermore, Table. 1 shows that using the additional label embedding can improve the solution by 4.86 in terms of FID. This verifies that the additional label embedding is beneficial for improving model generalization.
4 Conclusion
In this technical report, we present a solution for 3D object detection of ICCV 2023 OmniObject3D Challenge. The key insight of our solution is to learn effective NeRFs and SDFs Representations with 3D generative adversarial networks for 3D object generation. Overall, this solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.

### v2

# Learning Effective NeRFs and SDFs Representations with 3D GANs for Object Generation

@@ Abstract
We present a solution for 3D object generation of ICCV 2023 OmniObject3D Challenge. In recent years, 3D object generation has made great process and achieved promising results, but it remains a challenging task due to the difficulty of generating complex, textured and high-fidelity results. To resolve this problem, we study learning effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Specifically, inspired by recent works, we use the efficient geometry-aware 3D GANs as the backbone incorporating with label embedding and color mapping, which enables to train the model on different taxonomies simultaneously. Then, through a decoder, we aggregate the resulting features to generate Neural Radiance Fields (NeRFs) based representations for rendering high-fidelity synthetic images. Meanwhile, we optimize Signed Distance Functions (SDFs) to effectively represent objects with 3D meshes. Besides, we observe that this model can be effectively trained with only a few images of each object from a variety of classes, instead of using a great number of images per object or training one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is among the top 3 in the ICCV 2023 OmniObject3D Challenge.
Index Terms:
3D Object Generation, Generative Adversarial Networks, NeRFs, SDFs
I Introduction
To devise a solution for 3D object generation in the ICCV 2023 OmniObject3D Challenge, we simultaneously consider effective 3D representation learning and advanced generation strategy. In terms of 3D representation learning, textured mesh [12] and Neural Radiance Fields (NeRFs) [13] embrace the great potential to effectively encode 3D object features. In particular, textured mesh representations can support individual texture map generation, which allows us to easily replace surface textures, while NeRFs-based representations excel at rendering high-fidelity images. Furthermore, when considering effective shape representation, there are a variety of options, ranging from explicit representations like point-based [14], mesh-based [15] and voxel-based [16] approaches to implicit ones such as the Signed Distance Functions (SDFs) [17, 18] and occupancy function [19]. In light of these cutting-edge techniques, we mainly focus on exploring effective NeRFs representations for 3D object generation while still optimizing SDFs representations to effectively represent objects with 3D meshes.
@@ I Introduction
Moving on to advanced generation strategy, there are usually different options, such as latent generation [11] and direct generation [20]. In particular, latent generation methods aim to generate the representation of the desired output in the latent space and decode the latent code back to the 3D space, while direct generation focus on directly generating the desired 3D representation without utilizing the latent space. Besides, from the perspective of generation process, generation strategies can be categorized into Generative Adversarial Networks (GANs) [9, 10], Variational AutoEncoder (VAE) [21], flow-based [22], Diffusion Models (DMs) [11], etc.. In our solution, we employ the efficient geometry-aware 3D GANs [10] as the backbone for 3D object generation. The overall framework of our solution is depicted in Fig. 1. Specifically, we aim to learn effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Inspired by the recent success of efficient geometry-aware 3D GANs, i.e., EG3D [10, 7], we employ it as the backbone and incorporate label embedding and color mapping [7] to enable model training on different taxonomies simultaneously. Next, we generate NeRFs-based representations for rendering high-fidelity synthetic images and SDFs to effectively represent objects with 3D meshes through a decoder. In addition, we empirically find that we can effectively train this model with only a few images per object from various classes, rather than training with a great number of images per object or optimizing one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
Figure 2: Images rendered by our solution with label embedding.
II Related Work
3D assets [23, 1, 24] are crucial for robotics and computer vision [25] because they provide accurate spatial representations and realistic environments essential for training, simulation, and object recognition. Unfortunately, the collection of 3D assets is time-consuming and labor-intensive [26, 1]. For example, when constructing the large-scale dataset, OmniObject3D [1], the objects are placed in a calibrated environment [1, 26] built using fiducial markers [27, 28, 29], and the 3D scanners sample the objects from different perspectives. These fiducial markers, though low-cost [30, 31, 32, 33, 34], require additional labor to deploy in the environment. To acquire high-quality scanning results and avoid motion blur caused by aggressive sensor movement [35], it takes 15 minutes to one hour to finish the sampling of an object, depending on the geometric complexity of the object. Generating 3D assets [9, 10, 7, 36] is of great significance because it allows us to learn from limited real-world data with constrained renderings, enabling the creation of more diverse, photorealistic, and scalable 3D content for applications in robotics, computer vision, and beyond.
The field of 3D reconstruction and generation has seen significant advancements in recent years, driven by the development of neural rendering techniques. Neural Radiance Fields (NeRF) [13] pioneered the use of coordinate-based neural networks to represent 3D scenes as continuous volumetric functions, enabling high-quality novel view synthesis. However, NeRF’s reliance on dense input views and its computationally intensive training process limit its scalability. To address these issues, NeuS [37] introduced a signed distance function (SDF)-based [17, 18] representation, which improves surface reconstruction accuracy by explicitly modeling the geometry of objects. This approach achieves better performance in scenarios with sparse input views and complex topologies.
Building on these advancements, recent works have explored efficient and compact representations for 3D data. Among these, the triplane latent representation proposed in [10] is a milestone for 3D generation, where 3D shapes are encoded into three orthogonal feature planes. This representation strikes a balance between expressiveness and computational efficiency, enabling faster training and inference while maintaining high fidelity.
Leveraging the triplane latent representation, researchers have integrated it with generative models to achieve promising results in 3D synthesis. GAN-based methods [7, 38] utilize adversarial training to generate realistic 3D shapes by learning the distribution of triplane codes from large-scale datasets. On the other hand, DMs [39, 11] have recently gained traction for their ability to generate highly detailed and coherent 3D structures. By iteratively denoising triplane codes, diffusion-based approaches achieve superior control over the generation process, enabling applications such as text/image-to-3D.
Figure 3: Images rendered by our solution without label embedding.
III Method
III-A Motivation of Using GANs instead of DMs
@@ III-B Solution Pipeline
III-C Model Training
Since we are simultaneously optimizing NeRFs and SDFs representations, we ensure the generator can initially generate a unit sphere by initializing our model with SDF-pretraining as [7]. Then, the generator is optimized using the Adam optimizer with a learning rate of 0.0025, while the discriminator [10] uses a learning rate of 0.002. The batch size is set to 32 and the model is trained for 300,000 iterations. Additionally, we also use adaptive discriminator augmentation [40] to improve model generalization. Overall, the model training objective is:
@@ III-C Model Training
IV Experiments
IV-A Dataset
@@
IV-B Efficient Model Training
Instead of training one model for each class, we train a single model with various types of objects. The model takes each image along with its camera pose, focal length, and object type. Besides, we empirically find that we can effectively train this model with only a few images per object from various classes, so we select the first 8 images of each object for model training. This improves the model training efficiency.
TABLE I: Improvement in terms of FID of our solution after using the additional label embedding.
@@ IV-B Efficient Model Training
NFI (CVPR’23) [7] 62.426
Ours w/o label embedding 62.335
Ours w/ label embedding 57.624
IV-C Results
Fig. 2 shows some visualization results of synthesis images generated by our solution on OmniObject3D. It can be seen that our approach is capable of generating high-fidelity images from different views for 3D object generation. Besides, Fig. 3 shows results of our solution without the additional label embedding. By comparing Fig. 2 with Fig. 3 and the results in Table I, we can see that without the additional label embedding, the quality of the synthesized images is significantly worse than that of the solution with label embedding. This verifies that the additional label embedding is beneficial for improving model generalization and performance. Furthermore, we compare our method with a state-of-the-art approach proposed in CVPR 2023, nerf-from-image (NFI) [7], as shown in Table I, in terms of FID.
IV-D Discussion
Our solution is one of the follow-up works of EG3D [10] that leverages 3D GANs and triplane latent representation. On the other hand, the combination of DMs and triplane latent representation [11, 39] has gained significant popularity in recent years. In fact, after the announcement of the winners of the ICCV 2023 OmniObject3D Challenge, we found that almost all the competitors used DM-based approaches. However, as introduced in Section II and Section III-A, the superiority of DMs lies in their stability during training and control during the generation process (i.e., conditional generation), while the OmniObject3D Challenge is an unconditional generation challenge. Thus, it is an interesting topic to explore how well 3D GANs can perform, especially when researchers tend to favor DMs as the preferred solution. From this perspective, our work provides some valuable and inspiring results, demonstrating that 3D GANs are still in the game.
V Conclusion
In this paper, we present a solution for 3D object generation in the ICCV 2023 OmniObject3D Challenge. The key insight of our approach is to learn effective Neural Radiance Fields (NeRFs) and Signed Distance Functions (SDFs) representations using 3D generative adversarial networks (GANs), enabling high-fidelity 3D object synthesis from multiple viewpoints. Our experimental results demonstrate that the proposed method effectively generates high-quality images with fine details and accurate geometry. We further find that incorporating label embeddings is highly beneficial for improving the model’s learning effectiveness, leading to better generalization and performance. Additionally, our approach achieves competitive results compared to a state-of-the-art method proposed at CVPR 2023, further validating its effectiveness. Overall, our solution ranks among the top three in the ICCV 2023 OmniObject3D Challenge, highlighting that the 3D GAN-based method is competitive in an era where diffusion models are becoming the dominant approach for 3D generation.